A certain company produces and sells frozen pizzas to public schools throughout the eastern United States. Using a very aggressive marketing strategy, they have been able to increase their annual revenue by approximately $10 million over the past 10 years. But increased competition has slowed their growth rate in the past few years. The annual revenue, in millions of dollars, for the previous 10 years is shown.

Year Revenue
1 8.43
2 10.94
3 12.88
4 14.11
5 16.41
6 17.11
7 18.47
8 18.55
9 18.50
10 18.33

B) Using Minitab or Excel, develop a quadratic trend equation that can be used to forecast revenue (in millions of dollars). (Round your numerical values to three decimal places.)

Tt = ____

C) Using the trend equation developed in part (b), forecast revenue (in millions of dollars) in year 11. (Round your answer to two decimal places.)

$ ____ million

Thank you!
Answer #1

B)

from regression from excel:

Tt =5.584+2.941t-0.166t2

C)

from above regression equation:

predicted value=5.584+2.941*11-0.166*11^2 = 17.85 million (please try 17.86 if this comes wrong due to rounding error)